# Laser pulse focusing and energetic electron generation by magnetized plasma lens

### v1

@@ IV Magnetic field induced focused electron beam generation
The high gradient magnetic field focuses the wave pulse very rapidly and thus a RCP wave pulse with minimum spot size is achieved at the plasma edge. The ECR condition then results in the dumping of the EM wave energy to the electrons very rapidly. This is consistent with one of our earlier studies Dhalia et al. (2024) where we have shown the conversion of EM field energy to electron kinetic energy at the ECR point. This has also been illustrated from figure 10. In subplot $(a)$ we depict the EMF energy as a function of time and the transverse dimension $y$. We have averaged it over the $x$ dimension here. It should be noted that it maximizes when the pulse gets focused inside the plasma. However, at later stage the EMF energy gets weakened. This happens as the EMF energy gets transferred to the electron kinetic energy at the resonance point. This interdependence gets clearly depicted from subplots $(b)$ and $(c)$. Here we have shown the EMF energy and the electron kinetic energy by color plots as a function of $x$ and $t$. It should be noted that the EMF energy gets weaker at the specific location $x=x_{ECR}$ which is the resonance point. Exactly at this location the electrons acquire a huge kinetic energy.
Figure 11 also shows the energy evolution with time. The graph shows that nearly $80\%$ of the pulse energy has been transferred to electrons. Particle trajectory of these randomly chosen 2000 electrons from focal spot location have been shown in figure 12(a). Figure (a) shows the trajectory with time on the color axis. The transverse dimension of the emerging beam is about $5l_{N}$. These ejected particles strongly follow the external magnetic field direction and divides into two electron beams and propagates away. Figure 12(b,c) shows, angular distribution of these particle at time $300\omega_{pe}^{-1}$ and $450\omega_{pe}^{-1}$ with color axis showing electron counts and radius of the circle represents energy in MeV. At $300\omega_{pe}^{-1}$ , these ejected electrons haven almost directional distribution and directed along the laser axis. Later in time they diverges and spreads along the perpendicular direction of laser axis due to the external magnetic field present in vacuum. Maximum electron energy about $E\sim 0.78m_{e}c^{2}=0.4MeV$ against the incident EM wave of normalized vector potential $a_{0}=0.08$.
@@
V Summary
The above discussion shows that when a linearly polarized electromagnetic wave pulse propagates through a cold, underdense, homogeneous plasma lens embedded in a slowly diverging type of external magnetic field , the spot size of wave pulse reduces to a minimum size through self-focusing. The electrons in the plasma under the influence of electromagnetic wave, drift along external magnetic field direction and causes a density compression and depression which modifies the refractive index of plasma the EM wave to focus along the propagation axis. Our analysis has shown that focusing is also dependent on shape of plasma. One of the main advantage of using underdense magnetic lens is that energy loss in plasma could be minimal and focused wavepulse can be obtained. The simulation results matches quite well with the analytical results of Jha et al. (2006, 2007). Self-focusing is strongly dependent on polarization, where RCP wave self-focus and LCP wave defocus. By increasing the gradient of magnetic field, a wave pulse can be focused inside the plasma at ECR resonance and can generate a strong collimated electron beam. The theory can find application in the focus of strong intense laser and microwave driven fusion and industrial plasma as well as particle accelerator schemes. Magnetic field induced focusing can play a major role in inertial confinement fusion as well.

### v2

@@ 4 Magnetic field induced focused electron beam generation
The high gradient magnetic field focuses the wave pulse very rapidly and thus a RCP wave pulse with minimum spot size is achieved at the plasma edge. The ECR condition then results in the dumping of the EM wave energy to the electrons very rapidly. This is consistent with one of our earlier studies (Dhalia et al., 2024) where we have shown the conversion of EM field energy to electron kinetic energy at the ECR point. This has also been illustrated from figure 11. In subplot $(a)$ we depict the EMF energy as a function of time and the transverse dimension $y$. We have averaged it over the $x$ dimension here. It should be noted that it maximizes when the pulse gets focused inside the plasma. However, at later stage the EMF energy gets weakened. This happens as the EMF energy gets transferred to the electron kinetic energy at the resonance point. This interdependence gets clearly depicted from subplots $(b)$ and $(c)$. Here we have shown the EMF energy and the electron kinetic energy by color plots as a function of $x$ and $t$. It should be noted that the EMF energy gets weaker at the specific location $x=x_{ECR}$ which is the resonance point. Exactly at this location the electrons acquire a huge kinetic energy.
Figure 12 also shows the energy evolution with time. The graph shows that nearly $80\%$ of the pulse energy has been transferred to electrons. Particle trajectory of these randomly chosen 2000 electrons from focal spot location have been shown in figure 13(a). Figure (a) shows the trajectory with time on the color axis. The transverse dimension of the emerging beam is about $5l_{N}$. These ejected particles strongly follow the external magnetic field direction and divides into two electron beams and propagates away.
Figure 13(b,c) shows, angular distribution of these particle at time $300\omega_{pe}^{-1}$ and $450\omega_{pe}^{-1}$ with color axis showing electron counts and radius of the circle represents energy in MeV. At $300\omega_{pe}^{-1}$ , these ejected electrons haven almost directional distribution and directed along the laser axis. Later in time they diverges and spreads along the perpendicular direction of laser axis due to the external magnetic field present in vacuum. Maximum electron energy about $E\sim 0.78m_{e}c^{2}=0.4MeV$ against the incident EM wave of normalized vector potential $a_{0}=0.08$.
@@
5 Summary
The above discussion shows that when a linearly polarized electromagnetic wave pulse propagates through a underdense homogeneous plasma lens embedded in a slowly diverging type of external magnetic field , it divides into RCP and LCP wave pulse and spot size of RCP wave pulse reduces to a minimum size through self-focusing. The electrons in the plasma under the influence of electromagnetic wave, drift along external magnetic field direction and causes a density compression and depression which modifies the refractive index of plasma and the EM wave focus along the propagation axis. Our analysis has shown that focusing is also dependent on shape of plasma. With just a $79\mu m$ thin shaped plasma lens $(c=0.005)$, magnetic field induced focusing reduces spot size of an incident RCP wave by almost $\sim 1/3.5$ and increases intensity by almost three times. One of the main advantages of using an underdense magnetic lens is that energy loss in plasma could be minimal and a focused wave pulse can be obtained. Further, it is also established that by increasing the gradient of the magnetic field, a wave pulse can be focused inside the plasma at ECR resonance and can generate a strong collimated electron beam. The theory can find applications in the focusing of strong intense lasers and industrial plasma as well as particle accelerator schemes. Magnetic field-induced focusing can play a major role in inertial confinement fusion as well.